A university dean wishes to estimate the average number of hours that freshmen study each week. The standard deviation from a previous study is a good indication that you can use 2.6 hours as the population standard deviation. How large a sample must be selected if the dean wants to be 93% confident of finding whether the true mean is within 45 minutes of the population mean? (6 points)
Answer #1

Solution

standard deviation =   =2.6h =156minutes

Margin of error = E = 45

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

sample size = n = [Z/2* / E] 2

n = ( 1.96* 156 / 45 )2

n =46.17

Sample size = n =47